OpenStudy (anonymous):

Translate this to english and say whether its true or false : āˆ€ X ∈ P(N), X āŠ† R

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

For every x is an element of the set P(N), x is a subset of the set R.

myininaya (myininaya):

That is how I read it.

myininaya (myininaya):

I'm not sure what set R represents and what set P(N) represents.

OpenStudy (anonymous):

well i don't know how to make the power set symbol on this.

OpenStudy (anonymous):

its powerset(N)

OpenStudy (anonymous):

R represents real number

myininaya (myininaya):

what does N represent

myininaya (myininaya):

naturals?

OpenStudy (anonymous):

Natural numbers yes

myininaya (myininaya):

So do you know what the powerset of the naturals look like?

myininaya (myininaya):

We will use the notation you defined here since I don't know how to make the powerset symbol either

myininaya (myininaya):

or the double back R or N

myininaya (myininaya):

\[N=\left\{ 1,2,3,4,... \right\}\]

OpenStudy (anonymous):

what is subset and what is powerset ? i never really understood that lol

myininaya (myininaya):

\[P(N)=\left\{ \emptyset , \left\{ 1 \right\} , \left\{ 1,2 \right\},\left\{ 1,2,3 \right\},..., N\right\}\]

myininaya (myininaya):

basically a powerset is a set that list all of the possible subsets

myininaya (myininaya):

so the list I just wrote out is all the possible subsets of the natural numbers

myininaya (myininaya):

P({1,2,3}) ={emptyset, {1},{2},{3},{1,2},{1,3},{2,3},{1,2,3} } You know you listed them all when you have listed 2^number of elements

myininaya (myininaya):

Like in that example there are 3 elements in {1,2,3} so there are 2^3 subsets of {1,2,3}

myininaya (myininaya):

Do you understand what a subset and a powerset is? And what the difference is?

myininaya (myininaya):

Like I just gave the powerset of {1,2,3} above all the sets listed in that set are subsets of {1,2,3} -- emptyset subset of {1,2,3} {1} subset of {1,2,3} {2} subset of {1,2,3} {3} subset of {1,2,3} {1,2} subset of {1,2,3} {1,3} subset of {1,2,3} {2,3} subset of {1,2,3} {1,2,3} subset of {1,2,3}

OpenStudy (anonymous):

is empty set basically just the number 0 ? thanks for your explanation btw im actually just taking it all in right now so sorry i havent responded lol

myininaya (myininaya):

{ } is the set containing nothing there are no members (not even the member 0

OpenStudy (anonymous):

hmm okay..

myininaya (myininaya):

{0} is the set containing 0

myininaya (myininaya):

but 0 is something

myininaya (myininaya):

not nothing

myininaya (myininaya):

i know that sounds weird

OpenStudy (anonymous):

right

OpenStudy (anonymous):

lol

myininaya (myininaya):

like you can say { } or emptyset or use that little symbol I used before

myininaya (myininaya):

\[\emptyset \]

myininaya (myininaya):

so basically the question is asking you if all the members of the powerset of the natural numbers is also a subset of the the real numbers

myininaya (myininaya):

the members of the powerset are the subsets of the natural numbers like the empty set is a member of the powerset of the natural numbers the {1} is a member of the powerset of the natural numbers {2} is a member of the powerset of the natural numbers {1,2,3,4,5} is a member of the powerset of the natural numbers N is a member of the powerset of the natural numbers

OpenStudy (anonymous):

so true ?

myininaya (myininaya):

I think I could do a better yep

myininaya (myininaya):

\[P(\mathbb{N} )=\left\{ \emptyset, \\ \left\{ 1 \right\} , \left\{ 2 \right\},...,\\ \left\{ 1,2 \right\} , \left\{ 1,3 \right\}, \left\{ 1,4 \right\},..., \\ \left\{ 1,2,3 \right\},...,\\ \mathbb{N} \\ \right\}\]

myininaya (myininaya):

It is impossible to list all of those sets but you get the point I think

OpenStudy (anonymous):

Yes lol, Thanks!!